OpenStudy (anonymous):

Will fan and Medal!!! These three points are collinear.True or False (3, 6), (-2, -9), (0, -4)

OpenStudy (anonymous):

eq. of any line through (3,6) and (-2,-9) is \[y-6=\frac{ -9-6 }{ -2-3 }\left( x-3 \right),y-6=3\left( x-3 \right)\] see if (0,-4) lies on it.